If 5.21 g of silver nitrate react with sodium phosphate what mass of silver phosphate is produced?
1
Expert's answer
2018-08-15T04:10:24-0400

Question:
If 5.21 g of silver nitrate react with sodium phosphate what mass of silver phosphate is produced?
m(AgNO3) = 5.21 g
m(Ag3PO4) -?

Solution:
3AgNO3 + Na3PO4 = Ag3PO4 + 3NaNO3
m(Ag3PO4) = x g
M(Ag3PO4) = 418.7 g/mol
M(AgNO3) = 169.9 g/mol

m(AgNO3)/3M(AgNO3) = x/M(Ag3PO4)
x = (5.21 * 418.7)/ (3*169.9) = 4.27 g

Answer: m(Ag3PO4) = 4.27 g